90 minutes 11/01/2010 Q1. (20 pts) Calculate the currents h, l2 , and I3 in each of the branches of the circuit V3-18V Vi-12V V2-3V
Answer #1

Current through 12 V, upward = i1
Current through 18 V, upward = i3
Current through 3 V, downward = i2
Using kirchoff's junction law
i1 + i3 = i2
Now using kirchoff's voltage law in right loop
18 + 3 = 5*i3 + 3*i2
21 = 5*i3 + 3*i2

Now using kirchoff's voltage law in left loop

12 - 3 = -3*i2 + 15*i1

3 = -i2 + 5*i1
Solving above darken three equation
i1 = 1.29 Amp
i3 = 2.14 Amp
i2 = 3.43 Amp
Please upvote.